# Adaptive Randomized Smoothing: Certifying Multi-Step Defences against Adversarial Examples

### v1

@@ 4.2 CelebA Benchmark: Classification Without Spatial Alignment
CelebA, $\sigma=0.12$ 94.7 (0.5) 95.7 (1.2) 96.3(0.9)
CelebA, $\sigma=0.75$ 88.7 (1.2) 89.0 (1.4) 92.3(0.9)
CelebA, $\sigma=1.5$ 77.7 (2.5) 74.3 (1.7) 81.0(2.2)
@@
To evaluate ARS on a more realistic task with natural spatial variation, we use the CelebA face dataset in its unaligned versio We focus on the “mouth slightly open” (label 21) binary classification task because mouth location and shape vary. Furthermore we hypothesize that the necessary input for this task is well-localized, which affords an opportunity for the masking model to reduce the effective input dimension. The dataset consists of images with varied resolution, and meta-data about the position of different features, including the mouth. To make a more challenging benchmark we randomly crop all images to $160\times 160$ pixels for more spatial variation. The only crop constraint is that the mouth is $\geq 10$ pixels from the edge to ensure sufficient input to solve the task. Figure 13 shows example images from the test set, their respective masks from ARS, and the baseline static mask.
@@
LABEL:fig:celeba-results shows the certified accuracy for ARS, Cohen et al. (2019), and static mask, for three levels of the noise $\sigma$. First, both baselines perform very similarly. We can see from Figure 13 that the static mask is approximately identity (notice the $\geq 0.99$ scale), with only very slight dimming on the edges. This is because the mouth is not centred in our benchmark, so there is no one-size-fits-all mask. Second, ARS is able to predict a sparse mask that focuses on areas likely to have the mouth. The mask adapts to each input at test time, which is what enables the sparsity without performance degradation. Third, this sparse mask leads to a large noise reduction, enabling ARS to drastically improve both standard and certified accuracy. For instance, with $\sigma=0.75$, ARS improves the accuracy from $88.7\%$ to $92.3\%$ (a $3.6$ points improvement), while the certified accuracy at $r^{\infty}=0.005$ jumps from $40.7\%$ to $71.3\%$ (a more than $30$ points improvement!). At lower noise ($\sigma=0.12$) there is no increase in accuracy, but certified accuracy increases from $81.7\%$ to $90.0\%$ at $r^{\infty}=0.001$). At larger noise ($\sigma=1.5$), the static mask struggles to perform as well, but ARS sees significant increases ($10$ percentage points in accuracy, and $20.0\%$ to $31.3\%$ in certified accuracy at $r^{\infty}=0.01$).
@@ 4.3 ImageNet Benchmark: Classification on the Standard Large-Scale Dataset
To evaluate the scalability of ARS we experiment on ImageNet with $\sigma=0.5$ (without any modification). We compare with Cohen et al. (2019), which we reproduce for this large-scale setting. We also compare with static mask, and random mask (our mask model randomly initialized without training) to evaluate the effect of blind maksing. We evaluate two versions of ARS: our regular model, trained end-to-end; and a version in which we fix the base classifier to that of Cohen et al. (2019) and only train our mask model for $10$ epochs (with the Adam optimizer and constant learning rate). Figure 14 shows the certified accuracy for all models. We make three observations. First, by only training a mask model ARS is able to improve results by $1$ percentage point (from $57\%$ to $58\%$). Certified accuracy remains close to that of Cohen et al. (2019) at all radius, but ARS can certify slighlty larger radius: At $r^{\infty}=0.005$, the certified accuracy of Cohen et al. (2019), static mask, and random mask are 00, while that of ARS is still $15\%$ (which drops to 00 shortly after).
Figure 14: ImageNet $\sigma=0.5$.
Second, ARS trained end-to-end reaches an accuracy of $60\%$, $3$ points higher than that of Cohen et al. (2019). The certified accuracy is higher by similar margins up to $r^{\infty}=0.003$, where it becomes a bit lower than that of Cohen et al. (2019), before again certifying a slightly better maximum radius, with $13\%$ accuracy at $r^{\infty}=0.005$. Third, a randomly initialized mask model yields an accuracy of $49\%$ showing that our ARS model benefits are not explained by random masking, which is consistent with the mask learning to reduce the input dimension while preserving input semantics important to the task. Overall, these results show that ARS is general in that its adaptivity does not over-specialize to our adaptation benchmarks, and that it can scale to large datasets and complex classification tasks.
@@ 5 Discussion
Limitations: The multi-step adaptivity of ARS improves certification at the cost of increased model size and computation for RS. This impacts both training and testing computation, and is especially costly in the context of RS due to the Monte-Carlo estimation of the model’s expected predictions (over several forward passes at inference time). While we empirically show improvement by ARS, it would be interesting and important to investigate how it combines with other RS improvements such as adversarial training (Salman et al., 2019), consistency regularization (Jeong and Shin, 2020), higher order certification (Mohapatra et al., 2020), double sampling (Li et al., 2022b), and denoising by diffusion (Carlini et al., 2022).
@@
To conclude: we introduced Adaptive Randomized Smoothing (ARS) to reconnect RS with DP theory, to propose a new two-step defence for deep image classification, and to rigorously analyze such adaptive defences that condition on inputs at test time. This framework opens promising avenues for designing ML models that are adaptively and soundly robust with provable guarantees about their updates on natural and adversarial inputs.
@@ Appendix B Proofs
Appendix C Mask Architecture
@@ Appendix D Experiment Details
train batch size 256 64 250
test batch size 20 100 25
@@
optimizer AdamW SGD AdamW
Mask lr 1e-3 5e-2 1e-3
Model weight decay 1e-4 - -
(UNet) momentum 0.9 - -
step size 30 100 40
gamma 0.5 1.0 0.5
@@
Table 3: Hyperparameters for training ARS.
Table 3 provides the details of our ARS models’ hyper-parameters.
@@
We also tune hyper-parameters for the Hong et al. [2022] baseline on our 20kBG benchmark, focusing on $\sigma=0.75$, at $k=48$. We perform grid search on $\beta$ (the parameter of the generalized normal distribution for the noise), and find that $\beta=2.25$ (close to a Gaussian, but with a wider more and shorted tails) performs best. We also tune other hyper-parameters. In all settings we use training batch size 400, 90 epochs, and Adam optimizer. For the $\sigma=0.75,0.5$ settings be use learning rate 0.001. For the $\sigma=0.12,1.5$ settings we use learning rate 0.0005 with a step learning rate scheduler (30 step size and $\gamma=0.2$).

### v3

@@ 4.2 CelebA Benchmark: Classification Without Spatial Alignment
CelebA, $\sigma=0.25$ 94.3 (0.5) 93.0 (0.8) 97.0 (0.8)
CelebA, $\sigma=0.5$ 91.0 (0.8) 91.7 (0.9) 94.3 (0.9)
CelebA, $\sigma=1.0$ 83.3 (0.9) 84.7 (2.6) 91.0 (1.6)
@@
To evaluate ARS on a more realistic task with natural spatial variation, we use the CelebA face dataset in its unaligned version. We focus on the “mouth slightly open” (label 21) binary classification task because mouth location and shape vary. The input part relevant to this task is likely well-localized, which affords an opportunity for the mask model to reduce the effective input dimension. The dataset consists of images with varied resolution, and meta-data about the position of different features, including the mouth. To create a challenging benchmark, we randomly crop all images to $160\times 160$ pixels, which creates spatial variation in the mouth’s position. The only crop constraint is that the mouth is $\geq 10$ pixels from the edge to ensure sufficient input to solve the task. Figure 10 shows example images from the test set, their respective masks from ARS, and the baseline static mask.
@@
LABEL:fig:celeba-results shows the certified accuracy for ARS, Cohen et al. (2019), and static mask, for three levels of the noise $\sigma$. First, both baselines perform very similarly. We can see from Figure 10 that the static mask is approximately identity (notice the $\geq 0.99$ scale), with only very slight dimming on the edges. This is because the mouth is not centred in our benchmark, so there is no one-size-fits-all mask. Second, ARS is able to predict a sparse mask that focuses on areas likely to have the mouth. The mask adapts to each input at test time, which is what enables the sparsity without performance degradation. Third, this sparse mask leads to a large noise reduction, enabling ARS to drastically improve both standard and certified accuracy. For instance, with $\sigma=0.5$, ARS improves the standard accuracy from $91.0\%$ to $94.3\%$ (a $3.3$ point improvement), while the certified accuracy at $r^{\infty}=0.004$ jumps from $45.0\%$ to $79.3\%$ (a more than $30$ point improvement!). At lower noise ($\sigma=0.25$) there is still an increase in standard accuracy from $94.3\%$ to $97.0\%$, and an increase in certified accuracy from $68.3\%$ to $84.7\%$ at $r^{\infty}=0.002$). At larger noise ($\sigma=1.0$), ARS sees significant increases ($7.7\%$ points in standard accuracy, and from $21.3\%$ to $49.3\%$ in certified accuracy at $r^{\infty}=0.008$).
@@ 4.3 ImageNet Benchmark: Classification on the Standard Large-Scale Dataset
To evaluate the scalability of ARS we experiment on ImageNet (without any modification) with $\sigma=0.25,0.5,1.0$. For each noise level, we compare with Cohen et al. (2019), which we reproduce for this large-scale setting. We evaluate two versions of ARS: our regular setting (End-To-End); and a version that fixes the base classifier to the model trained as in Cohen et al. (2019), and only trains our mask model for $10$ epochs (Pretrain). The certified accuracy is plotted in LABEL:fig:imagenet-end2end and the standard accuracy is reported in Table 3.
Setting/Approach Cohen et al. ARS (Pretrain) ARS (End-To-End)
ImageNet, $\sigma=0.25$ 66.5 (0.009) 67.4 (0.002) 65.7 (0.006)
ImageNet, $\sigma=0.5$ 57.2 (0.009) 56.0 (0.003) 57.4 (0.010)
ImageNet, $\sigma=1.0$ 43.6 (0.005) 43.8 (0.002) 44.5 (0.010)
Table 3: Standard test accuracy ($r=0$) on ImageNet. ARS maintains standard accuracy.
When only training the mask model, certified accuracy remains close to that of Cohen et al. (2019) at all radii and noise levels. ARS trained end-to-end improves both standard and certified accuracy. The standard accuracy (at $r=0$) increases from $57.2\%$ to $57.4\%$ and from $43.6\%$ to $44.5\%$ when $\sigma=0.5$ and $\sigma=1$, respectively. For $\sigma=0.25$, standard accuracy for ARS is close but slightly lower than Cohen et al. (2019), while the pretrained ARS outperforms Cohen et al. (2019) from $66.5\%$ to $67.4\%$. Appendix F discusses other ARS improvements without certification.
Certified accuracy increases at larger $\sigma$. For instance at $\sigma=0.5$, ARS improves certified accuracy at $r^{\infty}=0.001$ from $48.9\%$ to $50.5\%$. At larger noise $\sigma=1.0$, ARS improves certified accuracy at $r^{\infty}=0.005$ from $21.7\%$ to $23.1\%$. This shows that ARS’ adaptivity generalizes outside of the specialized benchmarks we designed, and can scale to large datasets and complex classification tasks.
@@ 5 Discussion
Limitations: The multi-step adaptivity of ARS improves certification at the cost of increased model size and computation for RS. This impacts both training and testing computation, and is especially costly in the context of RS due to the Monte Carlo estimation of the model’s expected predictions (over several forward passes at inference time). While we empirically show improvement by ARS, it would be interesting and important to investigate how it combines with other RS improvements such as adversarial training (Salman et al., 2019), consistency regularization (Jeong and Shin, 2020), higher order certification (Mohapatra et al., 2020), double sampling (Li et al., 2022b), and denoising by diffusion (Carlini et al., 2022). Lastly, our adaptive masking technique provides improved certificates for the $L_{\infty}$ norm, but does not have the same effect for other norms such as $L_{2}$ (see remarks in §2.4). It is plausible that ARS’ adaptivity can lead to improvements under alternative norms, by leveraging different DP mechanisms and updates. We leave this exploration for future work.
@@
To conclude: we introduced Adaptive Randomized Smoothing (ARS) to reconnect RS with DP theory, to propose a new two-step defence for deep image classification, and to rigorously analyze such adaptive defences that condition on inputs at test time. This framework opens promising avenues for designing models that are adaptively and soundly robust with provable guarantees about their updates on natural and adversarial inputs.
Acknowledgments and Disclosure of Funding
We thank Krishnamurthy Dvijotham and Ian Goodfellow for reviewing and providing feedback during the Google DeepMind publication process. We thank Motasem Alfarra for discussing variations on input-dependent variance for randomized smoothing. We are grateful for the support of the Natural Sciences and Engineering Research Council of Canada (NSERC) [reference number RGPIN-2022-04469], as well as a Google Research Scholar award. This research was enabled by computational support provided by the Digital Research Alliance of Canada (alliancecan.ca).
@@ Appendix B Proofs
Appendix C Experiment Details
C.1 Mask Architecture
@@
train batch size 256 64 300
test batch size 20 100 100
@@
optimizer AdamW SGD SGD
Mask lr 1e-3 5e-2 5e-2
Model weight decay 1e-4 - 1e-4
(UNet) momentum 0.9 - 0.9
step size 40 - 30
gamma 0.5 - 0.1
@@
Table 4: Hyperparameters for training ARS. Check Section C.3 for more details of CIFAR-10 hyperparameters.
C.2 Hyperparameter tuning for CelebA
Table 4 provides the details of our ARS models’ hyper-parameters. On the CelebA dataset, we tune the hyper-parameters in the ARS, Cohen et al. [2019], and static mask settings at $\sigma=0.75$. In all settings, we settle on SGD with learning rate 0.05 and a step learning rate scheduler (step size of 3 and $\gamma=0.8$) for the base classifier. In the static mask setting, we use SGD with learning rate 0.01.
C.3 Hyperparameter tuning for CIFAR-10 BG20k
For the base classifier $g$ in Cohen et al. [2019], Static Mask and ARS experiments, we tune the optimizer and its hyperparameters using $k=64,~{}\sigma=1.5$ as the testbed. Based on these tuning experiments, we chose AdamW as the optimizer with an initial learning rate of 0.01 and weight decay to 0.0001. We scale the learning rate by 0.1 every $30^{th}$ epoch, with a batch size of 256. For the mask model $w$ in ARS experiments, we again used the AdamW optimizer with an initial learning rate of 0.001, weight decay of 0.0001, whilst scaling the learning rate by 0.5 every $40^{th}$ epoch. We used the same hyperparameters for all $k$’s and $\sigma$’s for these setups.
For Súkeník et al. [2021], for $k=32,48,64$, we tune the hyperparameters using $k=64$ as the testbed. We kept the optimizer same as used in the author’s code (SGD), setting the initial learning rate to 0.01, momentum to 0.9 and weight decay to 0. We scaled the learning rate by 0.1 every $30^{th}$ epoch. We tune $k=96$’s parameters separately. We start with an initial learning rate 0.1, keeping the rest of hyperparameters same as for $k=32,48,64$. Note that for $k=96$, we could not get the standard accuracy to improve upon random baseline for $\sigma=\{0.5,0.75,1.0,1.5\}$, despite extensively tuning the learning rate.
$k$ $\sigma=0.12$ $\sigma=0.25$ $\sigma=0.5$ $\sigma=0.75$ $\sigma=1.0$ $\sigma=1.5$
32 2.25 2.25 2.25 2.25 2.25 2.25
48 2.0 2.0 2.25 2.0 2.25 2.0
64 2.25 2.0 2.25 2.0 2.0 2.0
96 2.0 2.0 2.25 2.25 2.25 2.25
Table 5: UniCR $\beta$ chosen for each $k$, $\sigma$ setting.
@@ C.3 Hyperparameter tuning for CIFAR-10 BG20k
For UniCR, we tune $\beta$ (the parameter of the generalized normal distribution for the noise) using $\sigma=0.75$ and $k=48$ as the testbed. We perform grid search on $\beta$ and find that $\beta=2.0$ or $\beta=2.25$ (Gaussian and close to a Gaussian, but with a wider more and shorted tails) perform best. For each $k,\sigma$ setting, we train 3 models with $\beta=2.25$ and $2.0$ and choose the $\beta$ giving highest mean standard accuracy. The chosen $\beta$ for each setting is given in Table 5. For each setting we also tune optimizer hyper-parameters. At $k=32$, we use SGD optimizer. We use a learning rate of 0.01, momentum of 0.9 and weight decay of 0.0005 with a step learning rate scheduler (30 step and $\gamma=0.1$). At $k=48,64,96$ we use training batch size 256, 100 epochs, and AdamW optimizer. We use a learning rate of 0.001 with step learning rate scheduler (30 step size and $\gamma=0.5$).
Appendix D Additional Results on CIFAR10 BG20k Benchmark
We show here full sweep of results for CIFAR-10 Bg20k benchmark for $k={32,48,64,96}$ and $\sigma={0.12,0.25,0.5,0.75,1.0,1.5}$ for all our baselines and ARS. Similar to results presented in Table 1, we report the mean accuracy and standard deviation over three seeds.
$k$ $\sigma$ Cohen et al. Static Mask UniCR Súkeník et al.Δ ARS Δ
$32$ $0.12$ 79 (0.7) 78.6 (0.9) 77.5 (1.) 71.8 (1.6) 78.5 (0.7)
$0.25$ 70.6 (1.) 73.9 (0.8) 69.8 (1.4) 68.6 (2.8) 72.6 (0.9)
$0.5$ 63.6 (2.) 64.8 (0.9) 62.8 (0.8) 59.1 (1.6) 64 (1.4)
$0.75$ 55.5 (0.8) 57.8 (0.8) 53.1 (0.8) 49.3 (0.2) 57.3 (0.6)
$1.0$ 48 (0.7) 47.3 (1.3) 46.1 (0.9) 44.6 (1.0) 49.3 (0.6)
$1.5$ 38.6 (0.2) 39.1 (0.8) 36.8 (0.8) 36.9 (0.1) 38.3 (1.0)
$48$ $0.12$ 80.1 (0.8) 80.9 (0.7) 77.0 (2.9) 73 (0.1) 83.6 (0.4)
$0.25$ 71.6 (1.0) 72 (2.0) 69.7 (0.8) 65 (0.07) 75.5 (1.0)
$0.5$ 64.3 (0.2) 64.1 (1.6) 60.3 (0.6) 53.5 (1.8) 66.0 (0.8)
$0.75$ 52.5 (1.2) 56 (0.8) 52.7 (0.6) 41.8 (2.4) 57.6 (1.5)
$1.0$ 42.5 (2.1) 45.1 (1.2) 44.3 (0.2) 34.1 (1.0) 47.6 (2.1)
$1.5$ 30.8 (0.2) 34.9 (2.4) 32.3 (0.2) 25.5 (1.0) 34.1 (2.9)
$64$ $0.12$ 80.9 (1.0) 81.3 (1.0) 76.7 (0.8) 73.8 (0.2) 82.3 (1.0)
$0.25$ 71.67 (0.9) 73.1 (3.2) 67.8 (0.5) 64.1 (0.8) 77 (1.8)
$0.5$ 61.6 (2.7) 64 (1.4) 58.7 (0.2) 45.1 (1.1) 65.4 (1.8)
$0.75$ 49.4 (2.) 51.6 (1.0) 50.7 (0.2) 31.3 (2.0) 56.5 (2.6)
$1.0$ 41.3 (1.8) 40 (0.5) 42.2 (0.6) 26.5 (0.7) 50.4 (2.5)
$1.5$ 28.3 (0.2) 25.9 (3.2) 30.0 (1.8) 15.8 (2.7) 30.6 (1.0)
$96$ $0.12$ 79.1 (0.8) 79.6 (0.4) 77.5 (1.1) 66 (0.7) 80.3 (3.7)
$0.25$ 65.3 (1.6) 71.8 (1.3) 68.8 (1.8) 45.5 (0.9) 78.3 (2.2)
$0.5$ 56.6 (2.4) 59.5 (1.4) 59.7 (1.3) 10.8 (2.3) 69.8 (1.2)
$0.75$ 45.3 (3.4) 48 (4.3) 50.2 (0.5) 10.2 (0.4) 63.6 (3.5)
$1.0$ 33.8 (3.8) 36.9 (0.5) 41.3 (2.4) 10.4 (0.3) 56.3 (2.3)
$1.5$ 26 (1.4) 25.1 (0.6) 31.0 (1.4) 10.8 (1.4) 39.1 (2.5)
D.1 Impact of Certification Test-set Size
[…]
$\sigma$ Samples Approach
Cohen et al. Cohen et al. w/ AdamW ARS
0.12 200 77.2 (2.7) 79.0 (0.7) 78.5 (0.7)
500 79.6 (2.1) 82.1 (0.4) 82.3 (1.1)
10k 80.3 (0.6) 82.7 (0.6) 83.4 (0.5)
0.25 200 70.7 (3.7) 70.6 (1.0) 72.6 (0.9)
500 72.4 (2.6) 75.6 (0.9) 75.9 (0.9)
10k 73.7 (2.6) 77.2 (0.2) 77.7 (0.1)
0.5 200 62.0 (0.4) 63.6 (2.0) 64.0 (1.4)
500 63.1 (0.7) 65.2 (0.9) 65.9 (0.7)
10k 64.3 (0.5) 66.1 (0.3) 67.9 (0.3)
1.0 200 45.0 (1.5) 48.0 (0.7) 49.3 (0.6)
500 45.9 (1.6) 49.1 (0.4) 50.5 (0.8)
10k 46.8 (1.7) 50.0 (0.5) 51.5 (0.4)
Table 7: $k=32$ standard accuracy ($r=0$) on CIFAR-10 (20kBG). We report the mean accuracy and standard deviation over three seeds. We compare three approaches: Cohen et al. [2019] trained with the hyper-parameters they report in their GitHub repository, Cohen et al. [2019] trained with the hyper-parameters we report in 4, and ARS. We train models with each approach at $\sigma=0.12,0.25,0.5,1.0$. The corresponding models are certified on 200, 500, and 10,000 samples.
D.2 CIFAR-10 BG20k figures
Figure 21, Figure 23, Figure 25 and Figure 27 shows figures of different stages in our ARS architecture (Figure 1). In all of these figures, $1^{st}$ row corresponds to input images $X$, $2^{nd}$ row corresponds to images right after $\mathcal{M}_{1}$, $3^{rd}$ row corresponds to ARS masks, $4^{th}$ row corresponds to element-wise product of input $X$ and ARS masks, $5^{th}$ row corresponds to images right after $\mathcal{M}_{2}$ and $6^{th}$ row corresponds to images right after averaging $\mathcal{M}_{1},\mathcal{M}_{2}$.
Figure 21: Figures at different stages in our ARS architecture for CIFAR-10 $k=32$ input images. Check Section D.2 for detailed information about each row
Figure 23: Figures at different stages in our ARS architecture for CIFAR-10 BG20k $k=48$ input images. Check Section D.2 for detailed information about each row
Figure 25: Figures at different stages in our ARS architecture for CIFAR-10 BG20k $k=64$ input images. Check Section D.2 for detailed information about each row
Figure 27: Figures at different stages in our ARS architecture for CIFAR-10 BG20k $k=96$ input images. Check Section D.2 for detailed information about each row
@@ Appendix E Additional Results on CelebA
$\sigma$ Cohen et al. Static Mask ARS
$0.12$ 94.7 (0.5) 95.7 (1.2) 96.3 (0.9)
$0.25$ 94.3 (0.5) 93.0 (0.8) 97.0 (0.8)
$0.5$ 91.0 (0.8) 91.7 (0.9) 94.3 (0.9)
$0.75$ 88.7 (1.2) 89.0 (1.4) 92.3 (0.9)
$1.0$ 83.3 (0.9) 84.7 (2.6) 91.0 (1.6)
$1.5$ 77.7 (2.5) 74.3 (1.7) 81.0 (2.2)
Table 8: Standard Accuracy ($r=0$) on CelebA (unaligned and cropped). We report the mean accuracy and standard deviation over three seeds.
We show here a full set of results for the CelebA benchmark over all $\sigma={0.12,0.25,0.5,0.75,1.0,1.5}$ for Cohen et al. [2019], static masking, and ARS. Similarly to the results in Table 2, we report the mean standard accuracy and standard deviation over three seeds in Table 8. The certified accuracies are plotted in LABEL:fig:celeba-full-appen. We see that at all $\sigma$, ARS has the highest standard accuracy.
Appendix F Additional Results on ImageNet
Figure 34: The localized ARS masks produce un-noised object regions after averaging. For $\sigma=1$.
Figure 35: For $\sigma=0.5$.
Figure 36: For $\sigma=0.25$.
Figures 34, 35 and 36 show how adaptive masking reduces the noise around areas that are important to classification for ImageNet. The images follow our architecture visualized Figure 1. The mask model is provided the first query noised images as input. The learned masks, presented in the bottom left, are sparse and concentrated around the area of interest (the bird, or any labelled object). The second query noised images after weighted average use the mask to clearly reduce the noise around the bird.
Setting/Approach Cohen et al. ARS (Pretrain) ARS (End-To-End)
ImageNet, $\sigma=0.25$ 68.5 (0.1) 69.5 (0.0) 70.1 (0.2)
ImageNet, $\sigma=0.5$ 60.7 (0.5) 63.2 (0.1) 64.2 (0.2)
ImageNet, $\sigma=1.0$ 47.9 (0.1) 52.1 (0.0) 53.8 (0.3)
Table 9: Test accuracy without certification on ImageNet.
An interesting observation is that on ImageNet, ARS yields larger improvements to the test accuracy without certification. We follow the certification procedure of Cohen et al. [2019], which first determines the predicted class, and then certifies it if and only if the probability of this predicted class is such that $\underline{p_{+}}\geq 0.5$ (that is, it groups all other classes into one non-predicted class). If this is not the case, the prediction will count as not certified at $r=0$, even if the predicted class is still correct. We keep this procedure for consistency with prior work, but on a task with a large number of classes like in ImageNet, the accuracy at $r=0$ can be much lower as the non-certified accuracy.
We noticed that ARS significantly improves this non certified accuracy, while not improving the accuracy at $r=0$ as much (see Section 4.3). In effect ARS leads to more correct predictions, but this (correct) predicted class has $\underline{p_{+}}<0.5$, so the accuracy at $r=0$ does not increase. Table 9 shows this effect by comparing the test accuracy (no certification at all) across all methods. Under ARS, the test accuracy increases from $68.5\%$ to $70.1\%$ when $\sigma=0.25$, from $60.7\%$ to $64.2\%$ when $\sigma=0.5$, and from $47.9\%$to $53.8\%$ for $\sigma=1.0$. In summary, ARS achieves best test accuracy (without certification) for all the noise levels. This suggests that ARS helps more than shown by the default certification approach, and that a finer analysis that accounts for the probability of all classes could yield further improvements in certified accuracy.